How much diluent must be added to a 2gram bottle of antibiotic powder to obtain a concentration of 125mg/5ml?
Answer #1

2 gram = 2*1000 = 2000 mg = (125*16) mg

If we want to dilute the powder to concentration of 125mg/5ml, we will have to add 16*5 = 80ml of diluent.